OpenStudy (anonymous):

I roll a fair dice 7 times. What is the probability that all the six sides of the dice come up? How should I calculate this?

OpenStudy (anonymous):

Why to the power of 7? It doesn't matter what the last throw will appear to be what value because as long as all 6 sides have shown up?

OpenStudy (callisto):

sorry i misunderstood the question

OpenStudy (callisto):

6! / 6^6 * 1 = 720/46656 = 5/324 i'm not good at probability

OpenStudy (anonymous):

oh I see... Thanks for your help! :)

OpenStudy (callisto):

glad to help :)

OpenStudy (anonymous):

By the way, should it be 7! or 6!?

OpenStudy (callisto):

6! i think

OpenStudy (callisto):

it's no. of choices

OpenStudy (anonymous):

I see. I wondered because there are 7 throws and it seems like it could permutate to have higher chances?

OpenStudy (callisto):

but i did missed sth...please also times 6 to the answer

OpenStudy (callisto):

ie 5/54

OpenStudy (anonymous):

why times another 6 to the answer?

OpenStudy (callisto):

ah... neglect the 6..

OpenStudy (callisto):

sorry :(

OpenStudy (anonymous):

haha...it's okay. Thanks. :)

OpenStudy (callisto):

welcome but you gotta check it.. i really did bad in probability!

OpenStudy (anonymous):

no problem. Thanks. :)